OpenStudy (anonymous):

find the amount of heat that flow in an hour through an aluminium rod,6 cm in diameter and 1.7 m long, if one end of the rod is mantained at a temperature of 120 c and the end at 30 c?(thermal conductivity of aluminium is 240 Wm-1 C)

OpenStudy (anonymous):

nurin,take a look at this link.i'm not sure the answer is accurate or not..but,try to check it out. http://malaysia.answers.yahoo.com/question/index?qid=20110822184508AAWecOl

OpenStudy (anonymous):

Q = mc(dT) However you are mixing up two distinct problems. The temperature change is 90º. The mass is volume x density. Density is around 2,700 kg m-^3. But you have given a U value. So if it is per metre the diameter doesn't matter. So it is just 1.7 x 240 x 3600. As a Watt is a Joule per second. So the heat energy will be 604,800 J

OpenStudy (vincent-lyon.fr):

By analogy with electricity, thermal resistance of the uniform rod is: \(R=\Large \frac {1}{conductivity}\frac{length}{area}\) Thermal resistance applies for steady state which is the case here. It links thermal power \(\Phi\) flowing through the rod and temperature difference \(\Delta T\) by: \(\Delta T=R\:\Phi\) Total heat transfer is \(Q=\Phi\times time\)